Please help me solve this. Thanks!
Suppose that $2000 is invested at a rate of 6% per year compounded continuously. What is the balance after 1 yr? After 2 yrs?

Answer by stanbon(75887)

Suppose that $2000 is invested at a rate of 6% per year compounded continuously. What is the balance after 1 yr? After 2 yrs?
------------------
A(t) = Pe^(rt)
A(1) = 2000*e^(0.06*1)
A(1) = 2000*e^0.6
A(1) = 2123.67
-----------------
A(2) = 2000e^(0.06*2)
A(2) = 2000e^0.12
A(2) = 2254.99
===============
Cheers,
Stan H.